# Modelling and Classifying the Components of a Literature Review

### v1

@@ 3.1 Task definition and Annotation Schema
In order to produce a characterisation of portions of text from scientific literature that can support systems for the development of more structured literature reviews, we propose categorising sentences from the related work sections into specific rhetorical types. We frame this as a single-label multi-class classification problem, where each sentence is assigned to the most appropriate category. The annotation schema presented in this paper includes seven categories: Overall, Research Gap, Description, Result, Limitation, Extension, and Other. These categories are defined in Table 1.
@@
The novel annotation schema was developed by building on the theoretical work of Khoo et al. [23], who proposed an influential coding schema for annotating the macro-level discourse structure of literature reviews. Although their schema is comprehensive, it was not designed with automatic systems in mind and therefore requires modifications to be applicable in practical settings. For example, some categories in the original schema partially overlap, and others have definitions that lack sufficient clarity, which poses challenges for consistent annotation and automated interpretation.
To address these issues, we followed the protocol of Khoo et al. and conducted an iterative annotation process on 22 research papers drawn from various disciplines, initially applying their original coding schema. We then systematically refined the schema by merging overlapping categories and splitting ambiguous ones, resulting in a new set of distinct, clearly defined classifications that are more amenable to reliable interpretation by AI systems.
The main change concerns the introduction of a clear distinction between i) sentences that discuss the overall research topic and ii) sentences that focus on individual studies. These two levels are not explicitly addressed in Khoo et al.’s original framework, but they are important to conceptualise, as this distinction clarifies the meaning of the categories and facilitates annotation both by human users and automated methods. For example, the categories what, description, and method in the original schema are applicable at both the topic level and the study level, but their function differs depending on context.
In our classification schema, the topic level is represented by the categories Overall, which provides a general overview of the research area, and Research Gap, which identifies unresolved issues or open questions in the field. The study level includes four categories specific to individual studies: Description, Result, Limitation, and Extension. We introduced the category Other to prevent forced, potentially inaccurate, category assignments by the model in situations of low confidence.
@@
Proposed schema Association Khoo et al. [23] schema
@@
Table 2 compares the coding schema proposed by Khoo et al. [23] with our revised schema. The primary differences lie in the redefinition of categories and the explicit separation between topic-level and study-level discourse. In particular, we deconstruct Khoo et al.’s broad meta-critique category, which may conflate critiques of either a topic or a specific study, into two categories: Research Gap, which signals the need for further research at the topic level, and Limitation, which identifies methodological or conceptual shortcomings in a particular study. Our schema also resolves the ambiguity in Khoo et al.’s interchangeable use of what and description across both levels. We use Overall exclusively for sentences that describe the research area as a whole, and reserve Description for sentences detailing individual studies. The Overall category also consolidates several of Khoo et al.’s categories (namely what, description, meta-summary, and brief-topics) into a single, more coherent label for topic-level summaries. The Extension category has also been reconceptualised. While Khoo et al. use current-study for sentences that refer to the current work in a general manner and can therefore be easily conflated with Description or Limitation, our definition captures the motivations underlying the current work. This includes the articulation of new ideas, the identification of contrasting perspectives, and the elaboration of existing approaches.
@@ 3.2 The Sci-Sentence Benchmark
To evaluate the ability of modern LLMs to classify sentences according to the annotation schema introduced earlier, we developed the Sci-Sentence Benchmark. Sci-Sentence includes 700 sentences manually annotated by domain experts, along with 2,240 automatically labelled sentences. These sentences were extracted from the introduction, related work, and limitations sections of 22 scientific papers, spanning a diverse range of disciplines, including Computer Science, Business, Education, Medicine, and Psychology.
Sci-Sentence was developed in three phases. First, we conducted a workshop involving domain experts to define the annotation guidelines and compute inter-annotator agreement on a sample of 140 sentences. Second, the same experts individually annotated an additional 560 sentences, resulting in a total of 700 manually annotated sentences. This process produced the first, fully manually annotated, version of Sci-Sentence, which included 560 sentences for training and validation, and 140 sentences for testing. Finally, we leveraged Sonnet 3.0 to generate a larger version of the training and validation set by producing four alternative versions of each of the 560 sentences. This approach enabled the use of a more extensive training dataset for automatic methods, while retaining the fully manually annotated test set for evaluation purposes.
In the following, we provide a more detailed account of the development process.
@@
To demonstrate the feasibility of the annotation task, the clarity of the category definitions, and the consistency of the expert annotations, we assessed inter-rater agreement. Specifically, we employed two metrics: Fleiss’ Kappa [85] and Gwet’s AC1 [86]. Fleiss’ Kappa was used to measure the overall agreement among the three raters across the entire annotation exercise [85]. In contrast, Gwet’s AC1 was applied to evaluate agreement at the category level. This choice was motivated by the fact that Gwet’s AC1 is designed to overcome certain limitations of Fleiss’ Kappa, providing a more robust and stable measure of inter-rater reliability under varying prevalence and marginal distribution conditions [86].
A Fleiss’ Kappa of 0.90 was achieved for the overall agreement among the three raters, indicating a high level of inter-annotator reliability [87]. Table 3 reports Gwet’s AC1 scores for category-specific agreement. In most cases, the agreement is above 0.80. The only exceptions are the Research Gap and Limitation categories, with agreement values of 0.78 and 0.75, respectively, which still represent substantial agreement according to established guidelines [87]. This strong general and category-specific agreement indicates that the annotation task is well-defined and that the experts were able to produce consistent labels. In turn, this suggests that the Sci-Sentence benchmark can serve as a high-quality resource for training downstream applications.
@@
Given the labour-intensive and time-consuming nature of the annotation process, we explored the use of semi-synthetic data. Unlike fully-synthetic samples, which are generated by mimicking the statistical properties of a dataset, semi-synthetic data refers to data generated considering representation of real-world objects, such as original sentences [88]. Specifically, we aimed to generate artificial sentences that replicate the characteristics of those found in the original dataset [89]. This approach has gained considerable momentum with the advent of LLMs [90]. The literature provides compelling evidence that semi-synthetic data can enhance dataset diversity [91], support threats detection in security [92], address missing values [93], mitigate algorithmic bias [94], and support privacy-preserving data sharing [95]. In line with recent studies [96, 97, 98], our experimental findings (see Section 5) confirm the effectiveness of this strategy.
To this end, we employed Sonnet 3.0 (accessed via Amazon Bedrock) to generate four semi-synthetic sentences for each original sentence in the training and validation sets only, leaving the test set unaltered [99]. We subsequently evaluated the generated sentences to verify that they were sufficiently syntactically distinct from their corresponding source sentences. For this purpose, we used the normalised Levenshtein distance (see Appendix A), which measures the similarity between two sentences on a scale from 0 (identical) to 1 (completely different). Specifically, we computed the normalised Levenshtein distance between each original sentence and its four new variants, as well as the average distance between each new sentence and the remaining three. If any of these distances fell below or equal to 0.20, indicating insufficient syntactic variation, these sentences were discarded and regenerated using the language model until they met the threshold.
@@ 4 Experimental Methodology
This section describes the experimental methodology related to the second objective. Specifically, we conducted a comprehensive evaluation of a broad set of state-of-the-art LLMs [100, 101] on the novel Sci-Sentence Benchmark, under both zero-shot learning (ZSL) and fine-tuning settings. We assessed 37 alternative solutions spanning a variety of model architectures, including encoder-only, decoder-only, and encoder-decoder, and covering a wide range of parameter sizes. We also tested both open-source and proprietary solutions.
@@ 4.1 Zero-Shot Learning Settings
Among the 12 models, 8 were open-source and varied in size from 2 billion to 123 billion parameters. The remaining 4 were proprietary models whose parameter counts are not publicly available but are widely believed to exceed 700 billion. To ensure a fair comparison, we applied the same prompt template across all models. This template, refined through iterative prompt engineering following best practices [102], consists of three components: 1) an objective description to help the language model understand the task; 2) a precise explanation of the seven classification categories; and 3) a detailed procedure that includes the sentence to be classified along with a clearly specified output format to facilitate automated parsing. For transparency and reproducibility, the full prompt template is provided in Appendix B.
@@ 4.2 Fine-tuning Settings
The models were fine-tuned using Google Colaboratory instances equipped with Nvidia V100 and L4 GPUs. For GPT-4o-mini, however, we relied on the OpenAI API. To mitigate the computational and memory constraints of Google Colaboratory, we adopted QLoRA [103], a method that quantizes model weights from high-precision (32-bit) to low-precision (8-bit) formats. This quantization significantly reduces both computational overhead and memory usage.
To further reduce the number of trainable parameters in our decoder models, we evaluated two optimisation techniques: Low-Rank Adaptation (LoRA) [104] and Noisy Embedding Instruction Fine-Tuning (NEFT) [105]. LoRA introduces small, trainable matrices derived from a low-rank decomposition of weight updates. During inference, these updates are combined with the original weights to produce the final output. In contrast, NEFT adds random noise to embedding vectors during training. We selected LoRA due to its widespread adoption as a standard optimisation method [106], and NEFT for its demonstrated effectiveness in improving performance [107, 108, 109, 110].
@@
In Section 5 (Results), we will focus on the best configuration for each model in terms of training data (base vs. augmented) and optimisation technique (NEFT vs. LoRA). However, the complete set of results is available in the associated repository [^7] Results in the Sci-Sentence Repository – https://github.com/fcobolanos/Classifying-the-Components-of-a-Literature-Review/tree/main/results.
@@ 4.3 Overview of the Models
With respect to model size, although the number of parameters is widely regarded in the literature as the standard metric, there is no clear consensus on the specific thresholds that distinguish small from large models [111]. For example, Liu et al. [112] define small models as those containing approximately one billion parameters, while Fu et al. [113] consider models with up to ten billion parameters to still fall within the small category. Due to this lack of agreement, we established our own thresholds to achieve a relatively balanced distribution of models across size categories.
@@ 5 Results
In this section, we report and discuss the results of our experiments. We begin with the ZSL experiments, followed by those involving fine-tuning. As mentioned in Section 4, all models were evaluated on the test set of the Sci-Sentence Benchmark. Model performance was assessed using Precision, Recall, and F1-score.
@@
5.1 Zero-Shot Learning Experiments
@@
5.2 Fine-tuning Experiments
@@
As discussed in Section 4.2, the fine-tuning experiments were performed using a range of configurations. In particular, each decoder was fine-tuned under four different settings, obtained by combining two training sets (base and augmented) with two optimisation methods (LoRA and NEFT). In contrast, encoders and encoder-decoder models were fine-tuned on both training sets using the standard fine-tuning procedure. To ensure clarity and avoid unnecessary detail, we report only the best-performing configuration for each model. A more detailed analysis of the effects of the training set and optimisation technique on performance is presented in Section 6.
Table 6 presents the F1-score, precision, and recall of the fine-tuned models, ranked by F1-score within each category.
@@
Notably, the Encoder category produced results comparable to the smaller decoder models. In particular, Sci-BERT—a BERT variant pre-trained on academic text and therefore well-suited for this task—achieved an F1-score of 92.8%, matching the performance of Gemma2-2B. This outcome is especially interesting because encoder models are generally faster and more scalable than small decoders. Thus, they offer an efficient solution for sentence classification with only a 1.5 percentage point drop in F1-score compared to the best-performing open model (SuperNova-Medius), and a 3.6 point drop compared to the top proprietary solution (GPT-4o-mini).
@@
To investigate how the different categories in the annotation schema influence the performance of various classifier solutions, we analyse the performance of the top models across categories. Table 7 reports the F1-score, precision, and recall of the best-performing models for each category type.

### v2

@@ 3.1 Task definition and Annotation Schema
To support automatic systems in creating more structured literature reviews, we will characterize text from the related work sections into specific rhetorical types. Specifically, we propose classifying each sentence in these sections into a single, specific rhetorical type. We frame this as a single-label multi-class classification problem, where each sentence is assigned to the most appropriate category. The annotation schema presented in this paper includes seven categories: Overall, Research Gap, Description, Result, Limitation, Extension, and Other. These categories are defined in Table 1.
@@
This annotation schema was developed by building on the theoretical work of Khoo et al., (2011), who proposed an influential coding schema for annotating the macro-level discourse structure of related work sections. However, the original schema was not optimised for automated processing and therefore requires revision to ensure its suitability for machine learning classifiers. Some categories have soft boundaries that partially overlap, and their definitions depend on substantial contextual understanding, which hinders both consistent human annotation and automated interpretation. In addition, several categories are overly broad, introducing potential ambiguity for downstream automatic systems that rely on precise classification. For example, the “what” category may refer either to a general concept in the field or to a specific methodology introduced in the same paper.
To address these issues, we followed the protocol of Khoo et al., (2011) and conducted an iterative annotation process on over 300 sentences extracted from research papers in Computer Science, Business, Medicine, and Psychology. We initially applied their original coding schema and then systematically refined it by merging overlapping categories and splitting ambiguous ones. This process resulted in a new set of distinct and clearly defined classifications that are more suitable for consistent interpretation by AI systems.
The main change concerns the introduction of a clear distinction between i) sentences that describe the overall research topic and ii) sentences that refer to individual studies. These two levels are not explicitly addressed in the original framework by Khoo et al., (2011), yet they are practically important for supporting both human and automated annotation. For example, the categories what, description, and method in the original schema can apply to either the topic level or the study level, and their function varies significantly depending on the context.
In our classification schema, each category is restricted to one of these two levels in order to avoid ambiguity. The topic level is represented by the categories Overall, which provides a general overview of the research area, and Research Gap, which highlights unresolved issues or open questions in the field. The study level includes four categories that are specific to individual studies: Description, Result, Limitation, and Extension.
We also introduced the category Other because an automatic classifier sometimes encounters sentences that do not match any predefined category. This additional category prevents the model from assigning an incorrect label when its confidence is low or when the text is unrelated to the existing categories.
@@
Proposed schema Association Khoo et al. Khoo et al., (2011) schema
@@
Table 2 compares the coding schema proposed by Khoo et al., (2011) with our revised schema. Notably, we deconstruct Khoo et al.’s broad meta-critique category, which may conflate critiques of either a topic or a specific study, into two categories: Research Gap, which signals the need for further research at the topic level, and Limitation, which identifies methodological or conceptual shortcomings in a particular study. Our schema also resolves the ambiguity in Khoo et al.’s interchangeable use of what and description across both levels. We use Overall exclusively for sentences that describe the research area as a whole, and reserve Description for sentences detailing individual studies. The Overall category also consolidates several of Khoo et al.’s categories (namely what, description, meta-summary, and brief-topics) into a single, more coherent label for topic-level summaries. The Extension category has also been reconceptualised. While Khoo et al. use current-study for sentences that refer to the current work in a general manner and can therefore be easily conflated with Description or Limitation, our definition captures the motivations underlying the current work. This includes the articulation of new ideas, the identification of contrasting perspectives, and the elaboration of existing approaches.
@@ 3.2 The Sci-Sentence Benchmark
To evaluate the ability of modern LLMs to classify sentences according to the annotation schema introduced earlier, we developed the Sci-Sentence Benchmark. Sci-Sentence includes 700 sentences manually annotated by domain experts, along with 2,240 automatically labelled sentences. These sentences were extracted from the introduction, related work, and limitations sections of scientific papers in Computer Science, Business, Education, Medicine, and Psychology.
Sci-Sentence was developed in three phases. First, we conducted a workshop involving domain experts to define the annotation guidelines and compute inter-annotator agreement on a sample of 140 sentences. Second, the same experts individually annotated an additional 560 sentences, resulting in a total of 700 manually annotated sentences. This process produced the first, fully manually annotated, version of Sci-Sentence, which included 560 sentences for training and validation, and 140 sentences for testing. Finally, we leveraged Sonnet 3.0 to generate a larger version of the training and validation set by producing four alternative versions of each of the 560 sentences. This approach enabled the use of a more extensive training dataset for automatic methods, while retaining the fully manually annotated test set for evaluation purposes. In the following, we provide a more detailed account of the development process.
@@
To demonstrate the feasibility of the annotation task, the clarity of the category definitions, and the consistency of the expert annotations, we assessed inter-rater agreement. Specifically, we employed two metrics: Fleiss’ Kappa (Fleiss, , 1971) and Gwet’s AC1 (Gwet, , 2008). Fleiss’ Kappa was used to measure the overall agreement among the three raters across the entire annotation exercise (Fleiss, , 1971). In contrast, Gwet’s AC1 was applied to evaluate agreement at the category level. This choice was motivated by the fact that Gwet’s AC1 is designed to overcome certain limitations of Fleiss’ Kappa, providing a more robust and stable measure of inter-rater reliability under varying prevalence and marginal distribution conditions (Gwet, , 2008).
A Fleiss’ Kappa of 0.90 was achieved for the overall agreement among the three raters, indicating a high level of inter-annotator reliability (Landis and Koch, , 1977). Table 3 reports Gwet’s AC1 scores for category-specific agreement. In most cases, the agreement is above 0.80. The only exceptions are the Research Gap and Limitation categories, with agreement values of 0.78 and 0.75, respectively, which still represent substantial agreement according to established guidelines (Landis and Koch, , 1977). This strong general and category-specific agreement indicates that the annotation task is well-defined and that the experts were able to produce consistent labels. In turn, this suggests that the Sci-Sentence benchmark can serve as a high-quality resource for training downstream applications.
@@
Given the labour-intensive and time-consuming nature of the annotation process, we also explored the use of semi-synthetic data. Unlike fully-synthetic samples, which are generated by mimicking the statistical properties of a dataset, semi-synthetic data refers to data generated considering the representation of real-world objects, such as original sentences (Joshi et al., , 2024). Specifically, we aimed to generate artificial sentences that replicate the characteristics of those found in the original dataset (Marwala et al., , 2023). This approach has gained considerable momentum with the advent of LLMs (Liu et al., 2024a, ). The literature provides compelling evidence that semi-synthetic data can enhance dataset diversity (He et al., , 2021), support threats detection in security (Myneni et al., , 2023), address missing values (Berti-Équille et al., , 2018), mitigate algorithmic bias (Li et al., 2022b, ), and support privacy-preserving data sharing (Yale et al., , 2019). In line with recent studies (Kaddour and Liu, , 2023; Li et al., 2023b, ; Li et al., 2024c, ), our experimental findings (see Section 5) confirm the effectiveness of this strategy.
To this end, we employed Sonnet 3.0 (accessed via Amazon Bedrock) to generate four semi-synthetic sentences for each original sentence in the training and validation sets only, leaving the test set unaltered (Stan et al., , 2025). We subsequently evaluated the generated sentences to verify that they were sufficiently syntactically distinct from their corresponding source sentences. For this purpose, we used the normalised Levenshtein distance (see Appendix A), which measures the similarity between two sentences on a scale from 0 (identical) to 1 (completely different). Specifically, we computed the normalised Levenshtein distance between each original sentence and its four new variants, as well as the average distance between each new sentence and the remaining three. If any of these distances fell below or equal to 0.20, indicating insufficient syntactic variation, these sentences were discarded and regenerated using the language model until they met the threshold.
@@ 4 Experimental Methodology
This section describes the experimental methodology related to the second objective. Specifically, we conducted a comprehensive evaluation of a broad set of state-of-the-art LLMs (Patel et al., , 2023; Kasneci et al., , 2023) on the novel Sci-Sentence Benchmark, under both zero-shot learning (ZSL) and fine-tuning settings. We assessed 37 alternative solutions spanning a variety of model architectures, including encoder-only, decoder-only, and encoder-decoder, and covering a wide range of parameter sizes. We also tested both open-source and proprietary solutions.
@@ 4.1 Zero-Shot Learning Settings
Among the 12 models, 8 were open-source and varied in size from 2 billion to 123 billion parameters. The remaining 4 were proprietary models whose parameter counts are not publicly available but are widely believed to exceed 700 billion. To ensure a fair comparison, we applied the same prompt template across all models. This template, refined through iterative prompt engineering following best practices (Berryman and Ziegler, , 2024), consists of three components: 1) an objective description to help the language model understand the task; 2) a precise explanation of the seven classification categories; and 3) a detailed procedure that includes the sentence to be classified along with a clearly specified output format to facilitate automated parsing. For transparency and reproducibility, the full prompt template is provided in Appendix B.
@@ 4.2 Fine-tuning Settings
The models were fine-tuned using Google Colaboratory instances equipped with Nvidia V100 and L4 GPUs. For GPT-4o-mini, however, we relied on the OpenAI API. To mitigate the computational and memory constraints of Google Colaboratory, we adopted QLoRA (Dettmers et al., , 2024), a method that quantizes model weights from high-precision (32-bit) to low-precision (8-bit) formats. This quantization significantly reduces both computational overhead and memory usage.
To further reduce the number of trainable parameters in our decoder models, we evaluated two optimisation techniques: Low-Rank Adaptation (LoRA) (Hu et al., , 2021) and Noisy Embedding Instruction Fine-Tuning (NEFT) (Jain et al., , 2023). LoRA introduces small, trainable matrices derived from a low-rank decomposition of weight updates. During inference, these updates are combined with the original weights to produce the final output. In contrast, NEFT adds random noise to embedding vectors during training. We selected LoRA due to its widespread adoption as a standard optimisation method (Mao et al., , 2024), and NEFT for its demonstrated effectiveness in improving performance (Zhao et al., 2024a, ; Laurençon et al., , 2024; Li et al., 2024e, ; Pradeep et al., , 2023).
@@
In Section 5 (Results), we will focus on the best configuration for each model in terms of training data (base vs. augmented) and optimisation technique (NEFT vs. LoRA). We report the complete set of results is available in the associated repository [^7] Results in the Sci-Sentence Repository – https://github.com/fcobolanos/Classifying-the-Components-of-a-Literature-Review/tree/main/results.
@@ 4.3 Overview of the Models
With respect to model size, although the number of parameters is widely regarded in the literature as the standard metric, there is no clear consensus on the specific thresholds that distinguish small from large models (Wang et al., 2024a, ). For example, Liu et al., 2024b define small models as those containing approximately one billion parameters, while Fu et al., (2023) consider models with up to ten billion parameters to still fall within the small category. Due to this lack of agreement, we established our own thresholds to achieve a relatively balanced distribution of models across size categories.
@@ 5 Results
In this section, we discuss the results of our experiments. As described in Section 4, all models were evaluated on the test set of the Sci-Sentence Benchmark, and their performance was assessed using precision, recall, and F1-score.
The first three subsections present an overall analysis of model performance. Section 5.1 reports the results obtained by the 12 ZSL decoder models. Section 5.2 describes the best-performing configurations for each of the 25 fine-tuned models. Finally, Section 5.3 compares the top-performing models across the seven categories of the schema. The following three subsections (Sections 5.45.6) provide an in-depth analysis to interpret and contextualise the results presented earlier. Specifically, Section 5.4 details the most prevalent model errors. Section 5.5 investigates the effects of applying optimisation techniques (LoRA, NEFT) to the decoder models. Finally, Section 5.6 evaluates the impact of data augmentation on model performance across different optimisation strategies and training data solutions (base vs. augmented).
@@
5.1 Zero-Shot Learning
@@
5.2 Fine-tuned Models
@@
As discussed in Section 4.2, the fine-tuning experiments were performed using a range of configurations. In particular, each decoder was fine-tuned under four different settings, obtained by combining two training sets (base and augmented) with two optimisation methods (LoRA and NEFT). In contrast, encoders and encoder-decoder models were fine-tuned on both training sets using the standard fine-tuning procedure. To ensure clarity and avoid unnecessary detail, we report only the best-performing configuration for each model. A more detailed analysis of the effects of the optimisation techniques and training sets on performance is presented in Section 5.5, and Section 5.6.
Table 6 presents the F1-score, precision, and recall of the fine-tuned models. Within each architectural type, the models are listed in order of decreasing F1-score. In this section, we discuss their average performance across all rethorical categories, while the next section examines the performance of the top models within each category.
@@
Notably, the Encoder category produced results comparable to the smaller decoder models. In particular, Sci-BERT, a BERT variant pre-trained on academic text and therefore well-suited for this task, achieved an F1-score of 92.8%, matching the performance of Gemma2-2B. This outcome is especially interesting because encoder models are generally faster and more scalable than small decoders. Thus, they offer an efficient solution for sentence classification with only a 1.5 percentage point drop in F1-score compared to the best-performing open model (SuperNova-Medius), and a 3.6 point drop compared to the top proprietary solution (GPT-4o-mini).
@@
5.3 Performance across rhetorical categories
To assess how the categories in the annotation schema influence the performance of different classifier architectures, Table 7 reports the F1-score, precision, and recall of the best models for each architecture type (encoder, encoder–decoder, and small, medium, and large decoder) across the seven categories of the annotation schema. The table also includes the performance of the ZSL decoders (D-ZSL).